OpenStudy (anonymous):

How do I figure out the following question. What is the simplified form of each expression? 7x^-8 * 6x^3 A. 42/x^5 B. 1/42x^5 C. 42x^11 D.13x^-5

OpenStudy (cggurumanjunath):

a

OpenStudy (anonymous):

\[(7x^{-8})(6x^3)\] \[(\frac{ 7 }{ x^8 })(6x^3)\] \[\frac{ 42x^3 }{ x^8 }\] \[42x ^{-5}\] \[=\frac{ 42 }{ x^5 }\]

OpenStudy (anonymous):

so a

OpenStudy (anonymous):

Thanks for showing your work :)

OpenStudy (anonymous):

you're welcome :)